A bakery finds that the price they can sell cakes is given by the function p = 580 − 10x where x is the number of cakes sold per day, and p is price. The total cost function of the company is given by c = (30+5x) 2 where x is previously defined and c is total cost.
Find the revenue and marginal revenue functions [Hint: revenue is price multiplied by quantity i.e. revenue = price × quantity] (3 marks)
1
Expert's answer
2020-05-06T16:36:14-0400

revenue = price × quantity

R = p×q.

In our case quantity is the number of cakes sold per day or x and price= 580-10x.

R = p×x,

R = (580 - 10x)x= 580x - 10x2,

marginal revenue MR:

MR = dR/dx = 580 - 20x.

Answer: R = 580x - 10x2; MR = 580 - 20x